# Online Process Reward Leanring for Agentic Reinforcement Learning

### v2

@@ Abstract
Large language models (LLMs) are increasingly trained with reinforcement learning (RL) as autonomous agents that reason and act over long horizons in interactive environments. However, sparse and sometimes unverifiable rewards make temporal credit assignment extremely challenging. Recent work attempts to integrate process supervision into agent learning but suffers from biased annotation, reward hacking, high-variance from overly fine-grained signals or failtures when state overlap is rare. We therefore introduce Online Process Reward Learning (OPRL), a general credit-assignment strategy for agentic RL that integrates seamlessly with standard on-policy algorithms without relying on additional rollouts or explicit step labels. In OPRL, we optimize an implicit process reward model (PRM) alternately with the agent’s policy to transform trajectory preferences into implicit step rewards through a trajectory-based DPO objective. These step rewards are then used to compute step-level advantages, which are combined with episode-level advantages from outcome rewards for policy update, creating a self-reinforcing loop. Theoretical findings guarantee that the learned step rewards are consistent with trajectory preferences and act as potential-based shaping rewards, providing bounded gradients to stabilize training. Empirically, we evaluate OPRL on three distinct agent benmarks, including WebShop and VisualSokoban, as well as open-ended social interactions with unverfiable rewards in SOTOPIA. Crucially, OPRL shows superior performance over frontier LLMs and strong RL baselines across domains, achieving state-of-the-art results with higher sample-efficiency and lower variance during training. Further analysis also demonstrates the efficient exploration by OPRL using fewer actions, underscoring its potential for agentic learning in real-world scenarios.
@@ 1 Introduction
LLMs are rapidly evolving from passive generators into autonomous agents that can reason, act, and adapt strategies over long horizons, including search and research agents (Jin et al., 2025; OpenAI, 2025a), mobile and web navigators (Furuta et al., 2024; Bai et al., 2024), software engineering assistants (Yang et al., 2025; Wei et al., 2025a), game players (Wang et al., 2025), and social or embodied intelligence (Liu et al., 2025; Lu et al., 2025). Unlike conventional RL for LLM post-training in static, single-turn tasks (Ouyang et al., 2022; Shao et al., 2024), training LLM agents in dynamic, interactive environments faces particular challenges: (1) rewards are typically sparse and delayed, complicating credit assignment to intermediate actions; (2) trajectories are long and non-Markovian at the token level, with each step consisting of a chain-of-thought (CoT) (Wei et al., 2022) and an executable action, inflating variance when credit is pushed to individual tokens; and (3) environments and counterparts are non-stationary, open-ended and often come with unveriafiable rewards (e.g, dialogues). Consequently, trajectory-level optimization with a single outcome reward (Wang et al., 2025; Chen et al., 2025; Wei et al., 2025b) suffers from temporal credit assignment, yielding high-variance policy learning, brittle exploration, and limited gains on agent tasks.
@@
Recent work has attempted to solve these problems particularly through process supervision in RL. For example, Zeng et al. (2025); Zou et al. (2025); Zhang et al. (2025b) provide denser feedback at intermediate steps but require human annotation or hand-crafted heuristics that are costly, biased, and vulnerable to reward hacking. Generative reward models (GRMs) (e.g, LLM-as-judge) (Liu et al., 2025; Zha et al., 2025) that predicts criticality or correctness for each step reduce annotation overhead but can be noisy and inconsistent across domains. Token-level process reward models (PRMs) trained from outcome labels help in single-turn tasks (Yuan et al., 2025; Cui et al., 2025), but the rewards they offer tend to be overly fine-grained for agentic learning, amplifying variance and destabilizing training as trajectories grow. Other approaches (Feng et al., 2025; Choudhury, 2025) compute step-wise advantages by grouping identical states, an assumption that fails in open-ended language environments where state overlap is rare. Together, these limitations raise a core question for agentic RL: How can we design a credit assignment strategy that is label-efficient and stable, scales to long-horizon, multi-turn interactions, and remains robust and generalizable to both verifiable and unverifiable rewards in open-ended environments?
@@
To address this, we propose Online Learning Process Rewards (OPRL), a general credit assignment strategy for LLM agents. OPRL trains a PRM alongside the policy using only on-policy trajectories and their outcome preferences. At each training step, the current policy generates rollouts, which are ranked by a rule-based verifier or an outcome reward model (ORM) to form positive–negative trajectory pairs. We then update the PRM with a DPO-derived objective on these pairs. The updated PRM induces an implicit reward for each action by measuring its relative preference over the previous policy snapshot. Since this reward is computed per turn, it provides dense feedback to guide exploration while staying coarse enough to keep variance under control. Policy optimization then combines two complementary advantages: an episode-level advantage from outcome rewards and a step-level advantage from the implicit step rewards, capturing both global task success and the contribution of individual actions. OPRL is compatible with standard on-policy RL algorithms, such as PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024), REINFORCE++ (Hu et al., 2025), without explicit step labels or extra rollout cost.
OPRL addresses the limitations of prior work along several dimensions: (1) Our method provides fine-grained, preference-consistent step credit without step labels by converting trajectory-level preferences into step-level guidance. Theoretical analysis shows these implicit step rewards are Bayes-consistent with Bradley–Terry trajectory preferences and implement potential-based shaping that preserves the optimal policy set (see Theorem 3.3); (2) OPRL stabilizes multi-turn RL training and reduces variance by optimizing at the turn level using implicit step rewards rather than per-token rewards. Specifically, maximizing the shaped objective by the policy performs a KL-descent step toward the PRM, yielding bounded gradients that stabilizes training. (see Proposition 3.4); (3) OPRL only relies on trajectory-level preferences, which can be sourced from either rule-based verifiers (e.g., success signals) or unverifiable ORMs (e.g., LLM judges) for open-ended environments with rarely overlapping states. This enables a unified credit assignment strategy across domains.
Experiments across three challenging agent benchmarks show that OPRL is effective, efficient, and robust to diverse multi-turn settings. In WebShop and VisualSokoban, it consistently outperforms closed-source frontier models and strong RL baselines, achieving state-of-the-art results with Qwen-2.5-7B(-VL) (Yang et al., 2024) as the base model. In SOTOPIA, an open-ended social interaction environment with unverifiable rewards, OPRL increases goal completion by up to 14% in self-chat and 48% when chatting with GPT-4o (OpenAI, 2024). Further analysis shows faster convergence with lower variance than outcome-only or token-level PRM baselines, indicating high sample efficiency and stable training. OPRL also boosts both episode- and step-level rewards while producing shorter episodes, suggesting more useful exploration with fewer unnecessary actions. Finally, ablations indicate that advantage-level fusion is critical, environment step penalties offer only modest gains, and token-level PRMs are too fine-grained for stable multi-turn training. Our contributions are threefold:
  • We propose OPRL, a universal fine-grained credit assignment method that converts trajectory-level preferences into dense step-level guidance for agentic RL.
  • We provide theoretical guarantees that the resulting implicit step rewards are preference-consistent and constitute potential-based reward shaping with bounded gradients to stabilize multi-turn RL training.
  • Empirical results and analysis demonstrate superior performance of OPRL over baselines in sample efficiency, robustness across RL algorithms, and generalization to open-ended environments with unverifiable rewards.
@@ 2 Preliminaries
We consider the LLM agent task as a multi-step decision-making process, where the agent interacts with the environment to achieve a long-term goal through sequential decisions given a task prompt $x\in p(X)$. At each timestep $t$, the agent receives an observation $o_{t}$ (e.g, adversary messages or environmental feedback) and responds with a textual action $a_{t}\in V^{L}$, where $V$ denotes the token vocabulary and $L$ is the maximum generation length [^1] Note that we prompt the LLM agent to produce a reasoning process before executing an action, and $a_{t}$ refers to the whole sequence consisting of both the reasoning and the action at each timestep.. The environment then returns a scalar reward $r_{t}$ and provides the next observation $o_{t+1}$. Until the last timestep $T$, the full episode consists of a trajectory $\tau=\{(o_{1},a_{1},r_{1}),...,(o_{T},a_{T},r_{T})\}$. However, in real-world scenarios (e.g., dialogues), rewards can be sparse or delayed, such as feedback provided only at the end of a trajectory or weak signals for intermediate steps. Consequently, it is challenging to assign credit to individual steps within a trajectory, especially when there are many interaction turns.
@@
RL addresses the agent task by optimizing the policy $\pi_{\theta}(a_{t}|o_{1:t},x)$ of an LLM, with the objective of maximizing the expected cumulative rewards during interactions. To estimate the gradient of the expected reward with respect to $\theta$, policy gradient methods are used, such as PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024), and REINFORCE++ (Hu et al., 2025). These RL algorithms mainly differ in the manner of estimating advantages for policy update. For example, PPO computes advantages with a learned value function using generalized advantage estimation (GAE). GRPO and RLOO are critic-free and form relative advantages within a batch of $N$ samples for the same prompt—GRPO centers (often normalizes) each reward by the group mean-while RLOO uses a leave-one-out mean. REINFORCE++ instead uses batch-normalized rewards as the baseline reward.
@@
Implicit rewards has shown effectiveness in reward modeling for LLM alignment by enabling models to infer reward structures without explicit labels. Besides preference learning (Rafailov et al., 2024b; Ethayarajh et al., 2024; Wu et al., 2025; Zhang et al., 2025a). implicit rewards have been also used as outcome rewards to evaluate the quality of model outputs (Hosseini et al., 2024; Zhong et al., 2025). Further, Rafailov et al. (2024a) demonstrates that DPO can automatically learn a Q-function. Beyond the use as ORMs or Q-functions, recent work explores the derivation of PRMs from implicit rewards for test-time reranking (Yuan et al., 2025) or single-turn RL training (Cui et al., 2025). In this setup, a PRM is trained with only outcome labels and parameterize the process reward for each token as $r_{\phi}(y_{t}):=\beta\log\frac{\pi_{\phi}(y_{t}|y_{<t})}{\pi_{\text{ref}}(y_{t}|y_{<t})}$, where $\pi_{\phi}$ represents the PRM and $\pi_{\text{ref}}$ the reference model. $y_{t}$ denotes the $t$-th token in the response $y$.
@@ 3 Method
In this section, we first provide an overview of our method with a definition of implicit step rewards for online RL. We then present theoretical analysis to justify that these step rewards offer faithful and stable guidance for policy learning.
Figure 1: Overview of OPRL. The LLM agent interacts with an environment or a user simulator to generate multi‑step rollouts. An ORM evaluates each trajectory and provides sparse outcome rewards. These outcome‑labeled rollouts are used to update a PRM, which induces implicit process reward for each step in the trajecory. Using both the implicit step rewards and the outcome rewards, the agent’s policy is updated through RL.
@@ 3.1 Overview
OPRL introduces an online PRM that is optimized alternately with the policy, transforming the tendency to prefer more optimal actions into a dense reward for each step, guiding fine-grained exploration and improvement of the agent. Figure 1 shows the overall training pipeline of our method: (1) sample on-policy rollouts using the current policy $\pi_{\theta}$; (2) construct preference pairs from the same on-policy data according to outcome rewards $r_{o}(\tau)$, and update the PRM for a training step via a dpo-derived objective defined in Eq. 2; (3) use the updated PRM to compute the implicit step reward defined in Eq. 1, and optimize $\pi_{\theta}$ for a training step according to Eq. 6. The alternating optimization of $\pi_{\phi}$ and $\pi_{\theta}$ creates a self-reinforcing loop, where the policy and the PRM iteratively enhance each other. Below we first provide a definition of the implicit step reward, and detail the training process for each model. Please refer to Appendix A for the detailed algorithm of OPRL.
Definition 3.1 (Implicit step rewards).
Let $\tau=(o_{1},a_{1},...,o_{T},a_{T})$ denote a trajectory produced by the current policy $\pi_{\theta}$. For action $a_{t}$ in the trajectory at step $t$, its implicit step reward is defined as:
@@
where $\pi_{\phi}$ represents the PRM that is updated online, $\pi_{\theta_{old}}$ is the previous snapshot of the policy, and $\beta\in[0,1]$ refers to a temperature that scales the reward. The implicit step reward measures how much more probable the current action is under the freshly learned PRM than under the old policy. Positive values indicate actions that $\pi_{\phi}$ believes to be responsible for recent improvements, while negative values highlight actions that should be discouraged.
Optimizing online PRM via DPO.
For scalable online PRM training, we simply optimize $\pi_{\phi}$ on positive-negative trajectory pairs sampled by the policy without inducing additional rollouts, and derive a DPO-style objective:
@@
where $\sigma$ is the logistic sigmoid and $\beta$ follows Eq. 1; $\tau^{+}$ is a positive trajectory that is preferred to the negative one $\tau^{-}$, both of which are labeled by outcome rewards [^2] In our experiments on WebShop and VisualSokoban, positive trajectories are those with success rates above 0, while for SOTOPIA, positive trajectories have goal completion score above 6.. See Theorem 3.3 for the proof that $r_{\phi}$ is an unbiased per-step reward estimator.
@@
We use GRPO as an example to illustrate how to incorporate implicit step rewards into policy training, despite that our method is compatible with any policy gradient algorithm. As shown in Figure 2, for each task prompt $x$, we sample $N$ trajectories $\{\tau_{1},...,\tau_{N}\}$ from the current policy $\pi_{\theta}$, and obtain its corresponding outcome rewards $\{r_{o}(\tau_{1}),...,r_{o}(\tau_{N})\}$ through a rule-based verifier or an ORM. We then compute the episode-level advantages for each sampled trajectory:
@@
where $R_{o}=\sum_{i=1}^{N}r_{o}(\tau_{i})$. Next, we use the latest PRM $\pi_{\phi}$ (from the previous training step) to obtain the implicit step reward for each action $a_{t}^{i}$ via Eq. 1, and compute the step-level advantages:
@@
where $R_{s}=\cup_{i,t}r_{\phi}(a_{t}^{i})$ denotes the whole set of step rewards in the $N$ trajectories to keep scales comparable across episode lengths. We then combine the two-level advantages for policy update:
@@
where $\alpha$ is a hyperparameter that balances the global correctness and local quality, and we set it to 1 in all our experiments. Finally, the policy is optimized using a surrogate objective that is widely used in Schulman et al. (2017); Shao et al. (2024); Ahmadian et al. (2024); Hu et al. (2025) but without KL-divergence penalty:
@@
where $\rho_{\theta}(a_{t}^{i})=\frac{\pi_{\theta}(a_{t}^{i}|\mathbf{o}_{t}^{i},\mathbf{x})}{\pi_{\theta_{\text{old}}}(a_{t}^{i}|\mathbf{o}_{t}^{i},\mathbf{x})}$ is the importance sampling ratio at the step-level, and $\epsilon$ is a hyperparamter that controls the clipping range of the importance (set to 0.2 in all our experiments).
Figure 2: The credit-assignment strategy of OPRL. In OPRL, the episode-level advantages $A^{E}(\tau)$ are computed using outcome rewards $r_{o}(\tau)$, while the step-level advantages $A^{S}(a)$ are calculated from implicit step rewards $r_{\phi}(a)$ learned online. The final advantages for policy update is a combination of these two-level advantages. Note that our method is not limited to group-based RL (sampling multiple trajectories per task prompt), despite that we show an example here.

### v3

# Agentic Reinforcement Learning with Implicit Step Rewards

@@ Abstract
Large language models (LLMs) are increasingly developed as autonomous agents using reinforcement learning (agentic RL) that reason and act in interactive environments. However, sparse and sometimes unverifiable rewards make it extremely challenging to assign credit when training LLM agents that serve as a policy. Recent work attempts to integrate process supervision into RL but suffers from biased annotation, reward hacking, high-variance from overly fine-grained rewards or failtures when state overlap is rare. We therefore introduce implicit step rewards for agentic RL (iStar), a general credit-assignment strategy that integrates seamlessly with standard RL algorithms without relying on additional rollouts or explicit step labels. Particularly, we alternatively optimize an implicit process reward model (PRM) with the policy model to generate implicit step rewards via a trajectory-based DPO objective. Theoretical analysis shows that this learning objective produces a step-wise reward function. Then the implicit step rewards are used to compute step-level advantages, which are combined with trajectory (or episode)-level advantages for policy updates, creating a self-reinforcing training loop. We evaluate our method on three challenging agent benchmarks, including WebShop and VisualSokoban, as well as open-ended social interactions with unverifiable rewards in SOTOPIA. Crucially, iStar shows superior performance over frontier LLMs and strong RL baselines across domains, achieving state-of-the-art results with higher sample-efficiency and training stability. Further analysis also demonstrates efficient exploration by iStar with increased rewards in both step- and episode-level while maintaining fewer steps to achieve task success. Code will be available soon.
@@ 1 Introduction
LLMs are rapidly evolving from passive generators into autonomous agents that can reason, act, and adapt strategies over long horizons, including search agents (Jin et al., 2025; OpenAI, 2025a), mobile and web navigators (Furuta et al., 2024; Bai et al., 2024), software engineering assistants (Yang et al., 2025; Wei et al., 2025a), and social or embodied intelligence (Liu et al., 2025; Lu et al., 2025). Unlike conventional RL for LLM post-training in static, single-turn tasks (Ouyang et al., 2022; Shao et al., 2024), training LLM agents in interactive environments faces particular challenges: (1) rewards are typically sparse and delayed, complicating credit assignment to intermediate actions; (2) trajectories are long and non-Markovian in token level, with each step consisting of a chain-of-thought (CoT) (Wei et al., 2022) and an executable action, inflating variance when credit is pushed to individual tokens; and (3) environments and counterparts are non-stationary, open-ended and often come with unverifiable rewards (e.g, dialogues). Consequently, trajectory-level optimization with a single outcome reward (Wang et al., 2025; Wei et al., 2025b) suffers from credit assignment, yielding high-variance policy learning, brittle exploration, and limited gains on agent tasks.
@@
Recent work has attempted to solve these problems particularly through process supervision in RL. For example, Zeng et al. (2025); Zou et al. (2025); Zhang et al. (2025b) provide denser feedback at intermediate steps but require handcrafted step labels that are costly, biased, and vulnerable to reward hacking. Generative reward models (GRMs) (e.g, LLM-as-judge) (Liu et al., 2025; Zha et al., 2025) that predict criticality or correctness for each step reduce annotation overhead but can be noisy and inconsistent across domains. Implicit PRMs(Yuan et al., 2025; Cui et al., 2025) help in single-turn tasks, but the token-level process rewards tend to be overly fine-grained in agent training, amplifying variance and destabilizing training as trajectories grow. Other approaches (Feng et al., 2025; Choudhury, 2025) compute step-level advantages by grouping identical states, an assumption that fails in open-ended language environments where state overlap is rare. Together, these limitations raise a core question for agentic RL: How can we design a credit-assignment strategy that is label-efficient and stable, scales to multi-turn interactions, and remains robust and generalizable to (un)verifiable rewards in open-ended environments?
@@
To address this, we propose implicit step rewards for agentic RL (iStar), a general credit-assignment strategy for LLM agents. iStar jointly trains an implicit PRM with the policy model (the LLM agent) using trajectories collected online. At each training step, the policy model generates rollouts, which are ranked by an outcome reward verifier (or model) to form positive–negative trajectory pairs. The implicit PRM is optimized via a trajectory-based DPO objective on these pairs. The updated PRM then generates implicit step rewards for each action by measuring its relative preference over the previous policy snapshot. Since this reward is computed per step, it provides dense feedback to guide exploration while staying coarse enough to keep variance under control. When training the policy model, we combine two complementary advantages: an episode-level advantage from outcome rewards and a step-level advantage from the implicit step rewards, capturing both global task success and the contribution of individual actions. iStar is compatible with various existing RL algorithms, such as GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024), and DAPO (Yu et al., 2025), without relying on annotated step labels or additional rollouts.
Our method addresses the limitations of prior work along several dimensions: (1) iStar provides step credit without annotated labels through implicit rewards derived from a trajectory-based DPO objective, which is guaranteed to be a step-wise reward function learned from trajectory preferences (see Section 3.2); (2) iStar stabilizes training with multi-turn RL by learning implicit rewards at the step level rather than the token-level (Cui et al., 2025); (3) iStar only relies on trajectory preferences that can be sourced from (un)verifiable outcome rewards even in open-ended environments, enabling unified credit assignment across different domains.
Experiments on three challenging agent benchmarks show that iStar achieves superior performance over frontier LLMs and strong RL baselines, achieving state-of-the-art results in WebShop and VisualSokoban. In SOTOPIA, open-ended social interactions with unverifiable rewards, iStar increases goal completion by up to 14% in self-chat and 48% when chatting with GPT-4o (OpenAI, 2024). Our method can also be plugged into different RL algorithms to boost their performance. Further analysis shows higher sample efficiency and training stability of our method than vanilla RL and token-level PRM (Cui et al., 2025) baselines, as well as efficient exploration with increased rewards in both step- and episode-level and fewer steps to achieve a goal.
@@ 2 Preliminaries
We consider the LLM agent task as a multi-step decision-making process, where the agent interacts with the environment to achieve a long-term goal through sequential decisions given a task prompt $x\in p(X)$. At each timestep $t$, the agent receives an observation $o_{t}$ from the environment and responds with a textual action $a_{t}\in V^{L}$, where $V$ denotes the token vocabulary and $L$ the maximum generation length [^1] Note that we prompt the LLM agent to produce a reasoning process before executing an action, and $a_{t}$ refers to the whole sequence consisting of both the reasoning and the action at each timestep.. The environment then returns a scalar reward $r_{t}$ and transitions to the next state. Until the last timestep $T$, the full episode consists of a trajectory $\tau=\{(o_{1},a_{1},r_{1}),...,(o_{T},a_{T},r_{T})\}$. However, in real-world scenarios, rewards can be sparse and delayed, such as feedback provided only at the end of a trajectory.
@@
RL addresses the agent task by optimizing the LLM agent $\pi_{\theta}(a_{t}|o_{1:t},x)$, with the objective of maximizing the expected cumulative rewards during multi-turn interactions. Policy gradient methods are usually used, such as PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), RLOO (Ahmadian et al., 2024), and REINFORCE++ (Hu et al., 2025). These RL algorithms mainly differ in the manner of estimating advantages for policy updates. For example, PPO computes advantages with a learned value function using generalized advantage estimation. GRPO and RLOO are critic-free and form relative advantages within a group of $N$ responses for the same input prompt. GRPO centers (and often normalizes) each reward by the group mean, while RLOO uses a leave-one-out mean. REINFORCE++ instead uses batch-normalized rewards as the baseline reward.
@@
Implicit rewards have shown effectiveness in reward modeling for LLM alignment by enabling models to infer rewards without explicit labels (Ethayarajh et al., 2024; Wu et al., 2025; Zhang et al., 2025a). The implicit rewards are used to evaluate the quality of model outputs, such as DPO (Rafailov et al., 2024b)). Further, Rafailov et al. (2024a) demonstrates that DPO can automatically learn a Q-function. Beyond the use as reward models or Q-functions, recent work uses implicit token-level rewards for test-time reranking (Yuan et al., 2025) or single-turn RL (Cui et al., 2025): $r_{\phi}(y_{t}):=\beta\log\frac{\pi_{\phi}(y_{t}|y_{<t})}{\pi_{\text{ref}}(y_{t}|y_{<t})}$, where $\pi_{\phi}$ represents the token-level reward model and $\pi_{\text{ref}}$ the reference model. $y_{t}$ denotes the $t$-th token in the response $y$.
@@ 3 Method
In this section, we first provide an overview of our method with a definition of implicit step rewards for agentic RL. We then present theoretical analysis to justify the learning objective of the implicit PRM in iStar produces a step-wise reward function.
Figure 1: Overview of iStar. At each training step, an LLM agent interacts with the environment to generate multi‑step rollouts ranked by an outcome reward verifier (or model) to construct positive-negative trajectory pairs. These pairs are used to train an implicit PRM via a trajectory-based DPO objective, which generates implicit step rewards for each action produced by the agent. Finally, calculate step-level advantages using the implicit step rewards and episode-level advantages using outcome rewards to optimize the LLM agent (policy model) through RL.
@@ 3.1 Overview
In iStar, there is an implicit PRM optimized alternately with the policy model (LLM agent), transforming the tendency to prefer more optimal actions into dense step rewards to guide exploration and improvement of the agent. Figure 1 shows the overall training pipeline of our method. The alternating optimization between the implicit PRM and the policy model creates a self-reinforcing training loop to iteratively enhance each other. Below we first provide a definition of the implicit step rewards, and detail the training process for the implicit PRM and the policy model, respectively. Please refer to Appendix A for the detailed algorithm of iStar.
Implicit step rewards.
Let $\tau=(o_{1},a_{1},...,o_{T},a_{T})$ denote a trajectory produced by the LLM agent using policy $\pi_{\theta}$. For action $a_{t}$ in the trajectory at step $t$, its implicit step reward is defined as
@@
where $\pi_{\phi}$ represents the implicit PRM, $\pi_{\theta_{\text{old}}}$ refers to the previous snapshot of the policy $\pi_{\theta}$, and $\beta\in[0,1]$ is a temperature that scales the reward. The implicit step reward measures how much more probable the current action is under the freshly learned PRM than under the old policy. Positive values indicate actions that $\pi_{\phi}$ believes to be responsible for recent improvements, while negative values highlight actions that should be discouraged.
Optimizing implicit PRM via trajectory-based DPO.
For scalable online RL, we train the implicit PRM $\pi_{\phi}$ on positive-negative trajectory pairs sampled by the policy $\pi_{\theta}$ and derive a trajectory-based DPO objective:
@@
where $\sigma$ is the logistic sigmoid, and $\pi_{\phi}$, $\pi_{\theta_{\text{old}}}$ as well as $\beta$ follows Eq. 1. $\tau^{+}$ is a positive trajectory that is preferred to the negative one $\tau^{-}$, both of which are labeled by an outcome reward verifier (or model)  [^2] In our experiments on WebShop and VisualSokoban, positive trajectories are those with success rates above 0, while for SOTOPIA, positive trajectories have goal completion score above 6.. Particularly, Eq. 2 has two main differences from the original DPO objective (Rafailov et al., 2024b): (1) we use the previous policy snapshot $\pi_{\theta_{\text{old}}}$ as the reference model, whose parameters alter during training, instead of the initial policy model that keeps frozen; (2) preferences are learned from pairwise trajectories instead of single-shot responses. See Section 3.2 for the justification that Eq. 2 equivalents to a Bradley-Terry (BT) model with a step-wise reward function.
Figure 2: The credit-assignment strategy of iStar. In our method, episode-level advantages $A^{E}(\tau)$ are computed using outcome rewards $r_{o}(\tau)$, while step-level advantages $A^{S}(a)$ are calculated based on implicit step rewards $r_{\phi}(a)$ produced by the implicit PRM. The final advantages for policy updates is a combination of these two-level advantages.
@@
We use GRPO (Shao et al., 2024) as an example to illustrate how to integrate our implicit step rewards into policy learning, despite that our method is compatible with a variety of RL algorithms (Ahmadian et al., 2024; Hu et al., 2025; Yu et al., 2025; Zheng et al., 2025). As shown in Figure 2, for each task prompt $x$, we sample a group of $N$ trajectories $\{\tau_{1},...,\tau_{N}\}$ from the current policy $\pi_{\theta}$, and obtain its corresponding outcome rewards $\{r_{o}(\tau_{1}),...,r_{o}(\tau_{N})\}$ through an outcome reward verifier (or model). We then compute episode-level advantages $A^{E}$ for the group of trajectories:
@@
where $R_{o}=\sum_{i=1}^{N}r_{o}(\tau_{i})$. Next, we use the latest implicit PRM $\pi_{\phi}$ (from the previous training step) to obtain implicit step reward for each action $a_{t}^{i}$ via Eq. 1, and compute step-level advantages:
@@
where $R_{s}=\cup_{i,t}r_{\phi}(a_{t}^{i})$ denotes the whole set of step rewards in the $N$ trajectories. Particularly, when we use a group of trajectories starting from the same initial state (task prompt), we can generate various counterfactual scenarios. This helps us calculate a more accurate and stable estimate of the state-value baselines, leading to a better measure of the advantage for $a_{t}^{i}$. On the other hand, in a single trajectory, actions happen in different states and are influenced by noise specific to the policy used by the agent, leading to advantages in high variance.
Given episode-level advantages and step-level advantages, we combine them for policy updates:
@@
where $\alpha$ is a hyperparameter that balances the two-level advantages. The final advantage can differentiate not only between favorable and unfavorable trajectories but also beneficial and detrimental steps within a group of trajectories from the same initial state, enabling more dense rewards for policy learning in long-horizons.
Ultimately, given the advantages $A(a_{t}^{i})$, the policy model $\pi_{\theta}$ is trained using a surrogate objective that is widely used in Schulman et al. (2017); Shao et al. (2024) but without KL-divergence penalty:
@@
where $\rho_{\theta}(a_{t}^{i})=\frac{\pi_{\theta}(a_{t}^{i}|\mathbf{o}_{t}^{i},\mathbf{x})}{\pi_{\theta_{\text{old}}}(a_{t}^{i}|\mathbf{o}_{t}^{i},\mathbf{x})}$ is the importance sampling ratio at the step-level, and $\epsilon$ is a hyperparamter that controls the clipping range of the importance. Particularly, the step-level importance sampling ratio aligns well with our implicit step rewards to ensure low-variance training noise on multi-step rollouts, similar to (Zheng et al., 2025).
Remarks.
Alternating optimization between the implicit PRM and the policy model establishes a training loop that enhances stability and accelerates convergence. In particular, using rollouts produced by the current policy to train both models keeps their data distributions roughly consistent, minimizing off-policy bias and covariate shift. This consistency keeps implicit step rewards calibrated to the agent’s behavior, yielding dense and low-variance credit signals and preventing over- or underfitting the implicit PRM to ensure stable training. The result is a self-reinforcing loop: improved policies yield better preference data, refining the implicit PRM, which in turn delivers more accurate implicit step rewards to guide the policy.